OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12627 — Bordeaux <> Marseille
Scheduled Maintenance Report for Network & Infrastructure
Completed
There's 8x10 down between Bordeaux Marseille from 16/02 to 6:00 GMT, on 2 different providers (incident on the optical transmission).

Update 3 (UTC/CET(CEST)): 16/02/2015 08:38/ 09:38

Engineers are already on site. Incident is localized on the highway. Due to nature of the incident, access to exact break location is restricted by local authorities.

Update(s):

Date: 2015-02-18 09:22:11 UTC
Network provider has confirmed that the incident is resolved with a permanent fix.


Date: 2015-02-17 13:14:16 UTC
Restoration works continue. The estimated time of repair is 19:00 UTC/20:00 CET

Date: 2015-02-17 11:47:36 UTC
Half of our connections remounted at 09:43 GMT.

We still need to remount 4x10G.

We're waiting for information from our second provider.

Date: 2015-02-17 08:51:32 UTC
Restoration works are already in progress. We are chasing the engineers for ETR.

Date: 2015-02-17 08:39:00 UTC
For security reasons, restoration works have stopped overnight and will not resume untill 0700 UTC. No police protection possible.

Date: 2015-02-16 15:31:11 UTC
Permission to access the break location has been granted by the local authorities and preparatory works are currently ongoing.

Date: 2015-02-16 14:18:09 UTC
The fault has been further explained as vandalism. It appears that a canal was intentionally set on fire causing damages to two separate cables. The fiber provider's field crew is working to run three new 100 meter long cables and install new splice enclosure to begin the required repairs. The fiber provider anticipates fiber splicing to commence at approximately 18:00 GMT.

Date: 2015-02-16 14:17:51 UTC
The fault is localised in Toulouse. The optic cables melted following an accident/fire.

Preparation of work is in progress. We hope to get services back up before this evening.

Date: 2015-02-16 09:33:41 UTC
There's little traffic on the POP, and the traffic is rerouting via Paris directly via other secure connections.
Posted Feb 16, 2015 - 09:30 UTC